1. Consider an 2D infinite square well that extends from 0 to a in the direction, and from 0 to b in the y direction. a) write down the time independent wave function (ar, y for an electron inside this 2D infinite square well in terms of na and ny, the quantum numbers corresponding to the and y part of the wave function, respectively. b) Calculate the energy of the first excited state. Assume that a E 0.5 nm and b 1 nam. Express your answer in eV. c) Calculate the lowest degenerate energy level, and give all combinations of nar and ny that yield this energy level. Use the values of a and b given in the par b). Express your answer in e d) Draw a contour plot of the probability density function P (z, y) for n 1, ny 2. Be sure to label your axis and use the correct dimensions. Also, label the peak (s) of POr, y) with the actual value in S.I. units.
Answer #1

(a) The time independent wave function will be

\psi(x,y) = A \sin\left(\frac{n_x\pi x}{a} \right ) \sin\left(\frac{n_y\pi y}{b} \right )

The constant A can be determined by the normalisation condition:

\int_{0}^{a}\int_{0}^{b}\left | \psi(x,y) \right |^2dxdy = 1

A^2\int_{0}^{a}\sin^2\left(\frac{n_x\pi x}{a} \right )dx\int_{0}^{b}\sin^2\left(\frac{n_y\pi y}{b} \right )dy = 1

A^2 \frac{a}{2}\frac{b}{2} = 1

A = \frac{2}{ab}

So

\psi(x,y) = \frac{2}{ab} \sin\left(\frac{n_x\pi x}{a} \right ) \sin\left(\frac{n_y\pi y}{b} \right )

(b) Energy of a state is given by

E = \frac{\hbar^2\pi^2}{2m}\left(\frac{n_x^2}{a^2}+\frac{n_y^2}{b^2} \right )

Substituting the following values in the above equation:

m = 9.1*10-31 kg

hbar = 1.054*10-34m2 kg/s

a = 0.5*10-9m, b = 1*10-9m

For first excited state, nx=1, ny=2

So energy of first excited state is

E = 4.8\times 10^{-19}\text{ J} = \frac{4.8\times 10^{-19} }{1.6\times 10^{-19}}\text{ eV} = 3.01 eV

(c) Putting in values of a and b in the expression for energy:

E = \frac{\hbar^2\pi^2}{2m}\left(\frac{n_x^2}{(0.5\times 10^{-9})^2}+\frac{n_y^2}{(1\times 10^{-9})^2} \right )

E = \frac{\hbar^2\pi^2}{2m}\left( 4n_x^2 + n_y^2 \right)\left(\frac{1}{(10^-9)^2} \right )

By taking different combinations of nx and ny we can see that the lowest degenerate states are obtained for

nx=1, ny=4   and    nx=2, ny=2

The energy of lowest degenerate state can be calculated by substituting the above values:

E = 1.2\times 10^{-18}\text{ J} = 1.2\times 10^{-18}/1.6\times 10^{-19} \text{ eV} = 7.5 \text{ eV}

(d) The wavefunction for nx=1, ny=2 is

\psi(x,y) = 4 \sin\left(\frac{\pi x}{0.5} \right ) \sin\left(\frac{2\pi y}{1} \right )

I've plotted it in ROOT. You can use any other software you like:

4 sin(3.14*2*x) sin(2 3.14 y) 0.8 0.6 0.4 y (nm) 0.2 0 80 0.1 0.2 0.0.4 0.5 x (nm)